A child is pulling block of mass M = 5 kg on a rough surface

with a force of F1 = 30 N at an angle of θ = 30° to the

horizontal axis. The friction force F2 = 10 N.


Find out magnitude of resultant force


Draw the resultant force.


1
Expert's answer
2020-08-11T08:25:07-0400

Given

Applied force "F_1=30 N"

Frictional force "F_2=10 N"

The resultant force is

"R=\sqrt{(F_1^2+F_2^2+2F_1F_2cos150^o)}=\sqrt{((30)^2+(10)^2+2(30)(10)cos150^o)}=480 N"

Diagram: